# Persistent Homology for Structural Characterization in Disordered Systems

### v3

@@ I Introduction
For instance, the radial distribution function (RDF) characterizes the global structure of a system by analyzing its average density distribution, which is obtained by averaging the local density of individual particles across the system [25]. Besides, in the context of feature engineering for machine learning (ML), Atom-centered symmetry functions (ACSF) [26, 27] and smooth overlap of atomic positions (SOAP) [28] characterize the local environment of individual particles using coordinate-independent functions and Gaussian smoothing combined with spherical harmonics expansion. For global characterization, the ACSF and SOAP vectors of individual particles are typically concatenated into a single vector or subjected to straightforward transformations, such as compression into vectors of uniform length. Moreover, order parameters [29, 30] are commonly used to quantify the degree of order and symmetry. For local characterization, the Steinhardt’s bond-orientational order parameter measures local structural symmetry using spherical harmonics [31], while the ten Wolde’s approach extends this by defining bonding criteria between particles and constructing bonding networks to better distinguish ordered and disordered environments [32]. For global characterization, the bond-orientational order parameters for all particles are typically averaged to quantify the global degree of order, providing an estimate of LRO formation [31, 32]. Additionally, the static structure factor (SSF) investigates multiscale order by mapping the average density distribution of particles into the frequency domain using Fourier transformations, with the ability to capture SRO and LRO by adjusting the wave vector [33]. Medium-range order (MRO) also plays a significant role in understanding the relationship between macroscopic and macroscopic properties of materials, particularly in phase behaviors [34, 35]. However, the extraction of MRO, such as local connectivity [36] or motifs [25, 37, 38], indeed still relies on averaging or statistical processing of local information of particles.
@@
Topological Data Analysis (TDA) provides a robust and effective tool for representing and analyzing point cloud data, and through persistent homology (PH), a method from algebraic topology [50], we can encode the topological information of point clouds into a vectorizable representation [51, 52, 53]. It is notable that PH has already seen successful applications, such as in studying the global topological characteristics of osmolyte molecular aggregation and hydrogen-bonding networks [54], integrating machine learning to predict the structure-energy relationships of molecular clusters [55], and employing weighted approaches to uncover local topological features and propose new quantitative tools [56].
@@
    Sec. III presents the computational results and provides analysis, including the outcomes and interpretations of both ML-based and non-ML approaches. It also explores phase transitions and their onset by tracking changes in various metrics along the simulated trajectories.
@@
    Sec. IV concludes the article and discusses the directions of future research.
@@ II.2.3 Identifying Rearrangements
II.3 Sample Preparation
In sample preparation, molecular dynamics (MD) simulations [75, 76] serve as the core technique. To begin with, we utilized the MD simulations to generate multiple trajectories. Then, we sampled these trajectories to obtain system-level samples (a single frame from a trajectory) and particle-level samples (the local environment of an individual particle within a single frame of the trajectory). Finally, using these samples, we constructed different datasets for both ML and non-ML tasks.
II.3.1 Protocols of MD Simulation
II.3.2 Group Division
We generated trajectories for three groups according to the protocols provided in Sec. II.3.1. Specifically, Group 1 and Group 2 used Protocol 1, while two trajectories of Group 3 followed Protocol 2.
Group 1 and 2 considered 20 corresponding $T_{\text{final}}$ values which are as follows: $T_{1}=0.5$, $T_{2}=0.53$, $T_{3}=0.55$, $T_{4}=0.57$, $T_{5}=0.6$, $T_{6}=0.63$, $T_{7}=0.65$, $T_{8}=0.67$, $T_{9}=0.7$, $T_{10}=0.73$, $T_{11}=0.75$, $T_{12}=0.77$, $T_{13}=0.8$, $T_{14}=0.83$, $T_{15}=0.85$, $T_{16}=0.88$, $T_{17}=0.9$, $T_{18}=0.93$, $T_{19}=0.95$, $T_{20}=0.97$, i.e, $\text{Traj}(T_{i}),i\in[1,20]\cap\mathbb{N}$. Group $1$ includes temperatures with odd indices, i.e, the trajectories corresponding to $T^{(1)}_{i}=T_{2i-1}$; conversely, Group $2$ includes temperatures with even indices, i.e, the trajectories corresponding to $T^{(2)}_{i}=T_{2i}$, where $i\in[1,10]\cap\mathbb{N}$. Group 3 considered two trajectories, with their corresponding final $T_{\text{final}}$ values being: $T_{\text{final}}=0.5$ and $T_{\text{final}}=1.0$. We denote these two trajectories as $\text{Traj}_{\text{KA}}^{(1)}$ and $\text{Traj}_{\text{KA}}^{(2)}$, respectively.
II.3.3 Sample Labelling
For system samples, we use MSD and $Q_{6}$ for phase labeling, denoting the label for trajectory $\text{Traj}(T)$ at time step $t$ as $y^{(T)}_{\text{entire},t}\in\{\text{liquid},\text{crystal},\text{amorphous}\}$. All data are listed in Table 4 of Appendix C for verification.
II.3.4 Dataset Construction
II.4 ML-based Approach
II.4.1 Task Description
Therefore, we conducted two classification tasks. Task 1 tests the local descriptor $\mathbb{I}^{(P)}_{p}$ (Eq. 2) for identifying particles that are prone to rearrangement under specific conditions, while Task 2 evaluates the global descriptor $\mathbb{I}^{(P)}_{\text{entire}}$ (Eq. 3) for classifying phases (liquid, crystal, and amorphous) across the entire system.
For Task 1, we train on Set 1 and test on Set 2. For Task 2, we train on Set 3 and test on Set 4.
II.4.2 PCA on Feature Matrix
The feature matrix $\mathbb{X}$ is the result of feature engineering for samples. Specifically, it is constructed by concatenating the vectors $\mathbb{I}^{(P)}_{p}$ and $\mathbb{I}^{(P)}_{\text{entire}}$ corresponding to each sample into a single matrix.
In all ML tasks in this article, we apply Principal Component Analysis (PCA) [84] on the feature matrix $\mathbb{X}$, retaining a variance proportion $z$. Firstly, we standardize $\mathbb{X}$ to $\mathbb{X}_{\text{std}}$, ensuring each feature has a mean of 0 and variance of 1 using the formula $\mathbb{X}_{\text{std}}=(\mathbb{X}-\mu)/\sigma$, where $\mu$ is the mean and $\sigma$ the standard deviation.
@@ II.4.2 PCA on Feature Matrix
Note that, in the testing stage, it is essential to apply the same PCA transformation used during model selection and training. Specifically, Set 2 must use the PCA projection derived from Set 1, and Set 4 must use that derived from Set 3. This approach is necessary for two essential reasons. Firstly, it maintains a consistent feature space between training and testing, as the model was trained using the PCA-defined space from the training sets. Secondly, it prevents data leakage because recalculating PCA on the test sets could unintentionally incorporate test data characteristics into the model. Consistent PCA transformations are therefore essential for an unbiased evaluation of model performance.
@@
II.4.3 ML Models
The task of ML is to train a mapping $\mathbb{y}=f(\mathbb{X})$, where $\mathbb{X}$ is the feature matrix mentioned in Sec. II.4.2, and $\mathbb{y}$ is the vector formed by concatenating the labels $y^{(T)}_{p,t}$ or $y^{(T)}_{\text{entire},t}$ for each sample, as described in Sec. II.3.3. In other words, the goal is to train a mapping from the feature matrix to the target properties.
This article uses two kinds of Support Vector Machines (SVM) [85] as $f$:
@@ II.4.3 ML Models
    $f(\mathbb{X})=\mathbb{X}\cdot\mathbb{w}+u$
@@
For consistency, we set the hyperparameter optimization as follows: 10-fold cross-validation for validation, Matthews Correlation Coefficient (MCC)) [86] and accuracy as evaluation metrics, and a Bayesian optimizer [87] with 200 steps using Expected Improvement per second plus (EIps) as the acquisition function. For the multi-class strategy (only for Task 2), we try the methods of One-vs-One and One-vs-All at the same time, and select the best-performing one.
II.4.4 Shapley Values
@@
II.4.5 Scalar Field Trained by Linear SVM
We defined a scalar field, Global Softness (detailed in Eq. 5), derived from the ML results, to quantify the overall trend of fluidity in the system over time.
Global Softness. We use a linear SVM to obtain the optimal hyperplane $\mathbb{X}\cdot\mathbb{w}+u-\mathbb{y}=\mathbb{0}$ in the input space. For Task 1, the directed distance from the $p$-th sample to the hyperplane, calculated as $f(\mathbb{X}_{(p)}(t))/||\mathbb{w}||_{2}$, is defined as its softness [74], following the method by Cubuk et al. [90, 74]. This directed distance can be interpreted as a measure of the confidence that the sample belongs to a particular class, i.e., hard or soft. In essence, this approach abstracts the evolution of the local environment of a particle into a temporal scalar field [74]. By summing the softness of each particle and taking the average, we obtain the overall fluidity trend of the system. We use the global softness of the entire system to measure the overall fluidity trend at time step $t$:
Note that Global Softness is not our original method but an application of the SVM classification hyperplane.
II.5 Non-ML Approach
@@
SI is applicable to both local and global characterizations. As shown in Fig. 3, the PD of a homogeneous LJ system typically displays a characteristic pattern: $H_{0}$ persistence points align along a vertical line on the left, while $H_{1}$ and $H_{2}$ persistence points form two distinct but adjacent clusters in the lower-right corner. The value of SI quantifies the clarity of the boundary between these $H_{1}$ and $H_{2}$ clusters, providing a measure of particle ordering within the analyzed region. The figure illustrates the distribution of these clusters and their boundary, with the SI calculated from the mean difference and combined standard deviation of the clusters.
@@
Notably, SI is a purely mathematical and non-ML metric derived from the descriptive statistics of persistent homology, providing a complementary perspective to traditional physical methods.
III Results
III.1 ML-based Approach
III.1.1 ML Results and Analysis
The feature matrix $\mathbb{X}$ extracted from Sets 1 to 4 contains 1,600 columns ($\mathfrak{M}=\mathfrak{N}=40$), which is excessive for ML tasks, increasing computational complexity and the risk of overfitting. To address this, we applied Principal Component Analysis (PCA) [84] to reduce dimensionality by retaining a certain percentages of the explained variance, denoted as $z$.
We trained our model using Support Vector Machines (SVM) [85], of which the desicion function is denoted as $\mathbb{y}=f(\mathbb{X})=\mathbb{X}\cdot\mathbb{w}+u$, testing both linear [85] and RBF kernels [91]. The ML results are presented in Table 1. The linear kernel slightly outperformed the RBF kernel in both tasks, but regardless of which kernel was used, the accuracy remained consistently high.

### v4

@@ I Introduction
For instance, the radial distribution function (RDF) characterizes the global structure of a system by analyzing its average density distribution, which is obtained by averaging the local density of individual particles across the system [25]. Besides, in the context of feature engineering for machine learning (ML), Atom-centered symmetry functions (ACSF) [26, 27] and smooth overlap of atomic positions (SOAP) [28] characterize the local environment of individual particles using coordinate-independent functions and Gaussian smoothing combined with spherical harmonics expansion. For global characterization, the ACSF and SOAP vectors of individual particles are typically concatenated into a single vector or subjected to straightforward transformations, such as compression into vectors of uniform length. Moreover, order parameters [29, 30] are commonly used to quantify the degree of order and symmetry. For local characterization, the Steinhardt’s bond-orientational order parameter measures local structural symmetry using spherical harmonics [31], while the ten Wolde’s approach extends this by defining bonding criteria between particles and constructing bonding networks to better distinguish ordered and disordered environments [32]. For global characterization, the bond-orientational order parameters for all particles are typically averaged to quantify the global degree of order, providing an estimate of LRO formation [31, 32]. Recently, we have noted that the introduction of the perspective of inversion symmetry and its breaking [33, 34, 35, 36] has brought new insights to the relevant field. Additionally, the static structure factor (SSF) investigates multiscale order by mapping the average density distribution of particles into the frequency domain using Fourier transformations, with the ability to capture SRO and LRO by adjusting the wave vector [37]. Medium-range order (MRO) also plays a significant role in understanding the relationship between macroscopic and macroscopic properties of materials, particularly in phase behaviors [38, 39]. However, the extraction of MRO, such as local connectivity [40] or motifs [25, 41, 42], indeed still relies on averaging or statistical processing of local information of particles.
@@
Topological Data Analysis (TDA) provides a robust and effective tool for representing and analyzing point cloud data, and through persistent homology (PH), a method from algebraic topology [54], we can encode the topological information of point clouds into a vectorizable representation [55, 56, 57]. It is notable that PH has already seen successful applications, such as in studying the global topological characteristics of osmolyte molecular aggregation and hydrogen-bonding networks [58], integrating machine learning to predict the structure-energy relationships of molecular clusters [59], and employing weighted approaches to uncover local topological features and propose new quantitative tools [60]. PH has also been employed to reveal SRO and MRO in silica glass and Cu-Zr metallic glass [61]. Additionally, the Persistence Weighted Gaussian Kernel (PWGK), developed based on PH, has demonstrated stability and efficiency in feature engineering for machine learning (ML) tasks [62]. Furthermore, PH-based ML potential descriptors enable efficient prediction of the physical properties of amorphous carbon without the need for hyperparameter tuning [63]. These works have greatly inspired us.
@@
    Sec. III introduces the methods of data generation, including the protocols of molecular dynamics (MD) simulations, the approach of group division, the method of sample labeling, and the process of dataset construction.
@@
    Sec. IV introduces the computational experiments, presents the results, and provides an analysis, including the outcomes and interpretations of both ML-based and non-ML approaches. Additionally, it explores phase transitions and their onset by tracking changes in various metrics along the simulated trajectories.
  4. 4.
    Sec. V concludes the article and discusses the directions of future research.
@@ II.2.3 Identifying Rearrangements
II.3 Machine Learning (ML) Techniques
II.3.1 Basic Concepts
Machine learning (ML) can be employed to investigate the relationship between data instances and target variables. In this article, ML specifically refers to interpretable discriminative classification tasks with handcrafted feature engineering.
Firstly, data instances are transformed into a feature matrix through feature engineering, denoted as $\mathbf{X}=g(\mathbf{x})$, where $\mathbf{x}$ is a vector representing all instances. Each instance is mapped to a corresponding feature vector, forming the feature matrix $\mathbf{X}$, where the $i$-th row represents the $i$-th instance and the $j$-th column corresponds to the $j$-th feature.
Based on this representation, ML is applied to fit a composite function $\mathbf{y}=f(g(\mathbf{x}))=f(\mathbf{X})$, where $\mathbf{y}$ is the target property vector, and $f$ is the ML model that captures the relationship between the feature matrix $\mathbf{X}$ and $\mathbf{y}$. Here, we choose support vector machines (SVM) [82] as $f$, not only because of their interpretability but also due to the convenience of constructing a scalar field based on the directed distance from samples to the decision hyperplane in the input space when using a linear kernel. This allows for a more intuitive characterization of the target properties of the samples.
II.3.2 Principal Component Analysis (PCA)
In our research, the feature matrix $\mathbb{X}$ is constructed by concatenating the vectors $\mathbb{I}^{(P)}_{p}$ (defined in Eq. 2) and $\mathbb{I}^{(P)}_{\text{entire}}$ (defined in Eq. 3) corresponding to each sample into a single matrix.
Typically, $\mathbf{X}$ contains a large number of columns corresponding to multiple predictors, making dimensionality reduction necessary. Principal Component Analysis (PCA) [83] is employed to achieve this by applying a linear transformation that generates a new set of uncorrelated principal components. These components are ranked according to their explained variance, enabling the selection of a smaller subset while preserving most of the original information. Since the importance of each principal component is determined by its associated eigenvalue, ranking the eigenvalues quantifies their relative contribution to the total variance in the data.
The steps for performing PCA on X with an explained variance proportion of $z$ are as follows:
Firstly, we standardize $\mathbb{X}$ to $\mathbb{X}_{\text{std}}$, ensuring each feature has a mean of 0 and variance of 1 using the formula $\mathbb{X}_{\text{std}}=(\mathbb{X}-\mu)/\sigma$, where $\mu$ is the mean and $\sigma$ the standard deviation.
@@ II.3.2 Principal Component Analysis (PCA)
Note that, in the testing stage, it is essential to apply the same PCA transformation used during model selection and training. Firstly, it maintains a consistent feature space between training and testing, as the model was trained using the PCA-defined space from the training sets. Secondly, it prevents data leakage because recalculating PCA on the test sets could unintentionally incorporate test data characteristics into the model. Consistent PCA transformations are therefore essential for an unbiased evaluation of model performance.
@@
II.3.3 ML Models
This article uses two kinds of Support Vector Machines (SVM) [82] as the ML model $f$:
@@ II.3.3 ML Models
    $f(\mathbb{X})=\mathbb{X}\cdot\mathbb{w}+u$ (5)
@@
In this research, we set the hyperparameter optimization as follows: 10-fold cross-validation for validation, Matthews Correlation Coefficient (MCC)) [84] and accuracy as evaluation metrics, and a Bayesian optimizer [85] with 200 steps using Expected Improvement per second plus (EIps) as the acquisition function. For the multi-class strategy (only for Task 2), we try the methods of One-vs-One and One-vs-All at the same time, and select the best-performing one.
II.3.4 Scalar Field Trained by Linear SVM
As shown in Fig. 3, when explicitly choosing a linear SVM, the training objective is to determine an optimal decision hyperplane that maximizes the margin between the two classes. According to Eq. 5, the function of this decision hyperplane is:
$\mathbb{X}\cdot\mathbb{w}+u-\mathbb{y}=0$ (6)
This hyperplane is determined by a subset of data points closest to it, known as support vectors, which are marked with circles on the dashed lines in Fig.3. The support vectors not only define the placement of the decision boundary (the solid line in Fig.3) but also establish two support vector hyperplanes (the symmetric dashed lines in Fig. 3). These hyperplanes are parallel to the decision boundary, forming the margins between the two classes and ensuring that all other data points remain outside this margin.
Figure 3: A schematic diagram of a linear SVM, illustrating the decision hyperplane, support vectors, support vector hyperplanes, positive and negative samples, and the directed distance from a certain sample to the decision hyperplane. For convenience, this figure only illustrates the idealized two-dimensional case without loss of generality.
Furthermore, the directed distance from a sample point to the decision hyperplane, which forms a scalar field, can be used to quantify the confidence of the sample belonging to a particular class, as illustrated by the deep yellow directed line segment in Fig. 3.
In Sec. II.2.3, we introduced a binary classification system for particles, using Eq.4 to distinguish between soft particles, which are prone to rearrangement, and hard particles, which are more resistant to it. When the sample consists of particles, the directed distance of a corresponding point from the classification hyperplane represents the confidence level in the classification of a particle as “soft” or “hard”. A positive value indicates a soft particle, while a negative value corresponds to a hard particle. The magnitude of this value quantifies the tendency of a particle to undergo rearrangement.
Following the method by Cubuk et al. [86, 81], we define the directed distance from the corresponding point of the $p$-th particle to the decision hyperplane at time $t$ as the Local Softness of the particle at time $t$:
$\mathcal{s}_{(p)}(t)=\frac{f(\mathbb{X}_{(p)}(t))}{||\mathbb{w}||_{2}}=\frac{\mathbb{X}_{(p)}(t)\cdot\mathbb{w}+u}{||\mathbb{w}||_{2}}$ (7)
In essence, this approach represents the evolution of the tendency for particle rearrangement as a temporal scalar field [81].
For a system with $\mathscr{N}$ particles, the overall fluidity trend of the system can be obtained by summing the Local Softness of all particles and taking the average. We define this averaged scalar field as Global Softness, which quantifies the overall fluidity trend of the system at time step $t$:
Note that Global Softness is not our original method but an application of the SVM decision hyperplane.
II.3.5 Shapley Values
@@
II.4 Descriptive Statistics on PH (non-ML approach)
@@
The process of increasing $\epsilon$ from 0 can be interpreted as the emergence and disappearance of topological features. As defined in Eq. 9, the subscript $j$ denotes the index of each topological feature within the homology group, corresponding to each birth-persistence pair in the barcode data. The Betti number $\beta_{i}^{\max}$ indicates the total number of topological features in $H_{i}$, meaning the total number of birth-persistence pairs in dimension $i$. The time step $t$ captures the evolution of the system at different time points. Ultimately, SI quantifies the separation between the point clusters corresponding to different homology classes in the PD by calculating the sum of mean differences of the $(b,d)$ pairs in $H_{1}$ and $H_{2}$ to the sum of their standard deviations.
@@
As shown in Fig. 4, the PD of a homogeneous Lennard-Jones (LJ) system, generated by Protocol 1, as defined in Sec. III.1, typically exhibits a characteristic pattern:
@@
As shown in Fig. 4, SI provides an intuitive visual interpretation. Specifically, the PD of a homogeneous Lennard-Jones (LJ) system, generated by Protocol 1, as defined in Sec. III.1, typically displays a characteristic pattern: $H_{0}$ persistence points align along a vertical line on the left, while $H_{1}$ and $H_{2}$ persistence points form two distinct but adjacent clusters in the lower-right corner. The value of SI quantifies the clarity of the boundary between these $H_{1}$ and $H_{2}$ clusters, providing a measure of particle ordering within the analyzed region. The figure illustrates the distribution of these clusters and their boundary, with the SI calculated from the mean difference and combined standard deviation of the clusters.
@@
Notably, SI is a purely mathematical and non-ML metric derived from the descriptive statistics of persistent homology, providing a complementary perspective to traditional physical methods. Besides, SI is applicable to both local and global characterizations, as the objects they describe are essentially point clouds.
III Dataset Generation
In our study, molecular dynamics (MD) simulations [89, 90] serve as the core technique. To begin with, we utilized the MD simulations to generate multiple trajectories. Then, we sampled these trajectories to obtain system-level samples (a single frame from a trajectory) and particle-level samples (the local environment of an individual particle within a single frame of the trajectory). Finally, using these samples, we constructed different datasets for both ML and non-ML tasks.
We provide a comprehensive summary in Table 1, detailing the molecular dynamics (MD) simulation protocols, system categories, and simulation dynamics. Additionally, we outline potential structural transitions observed in the trajectories, classification of groups, extracted datasets, corresponding state sets, their respective tasks, and the figures generated from these datasets for reference.
III.1 Protocols of MD Simulation
III.2 Group Division
We generated trajectories for three groups according to the protocols provided in Sec. III.1. Specifically, Group 1 and Group 2 used Protocol 1, while two trajectories of Group 3 followed Protocol 2.
Group 1 and 2 considered 20 corresponding $T_{\text{final}}$ values, denoted as $T_{1},T_{2},\dots,T_{20}$, which form a monotonically increasing sequence with equal intervals (see Table 5 in Appendix E). The trajectory corresponding to $T_{j}$ is denoted as $\text{Traj}(T_{j})$, where $j\in[1,20]\cap\mathbb{N}$. Group $1$ includes temperatures with odd indices, i.e, the trajectories corresponding to $T^{(1)}_{i}=T_{2i-1}$; conversely, Group $2$ includes temperatures with even indices, i.e, the trajectories corresponding to $T^{(2)}_{i}=T_{2i}$, where $i\in[1,10]\cap\mathbb{N}$. Group 3 considered two trajectories, with their corresponding final $T_{\text{final}}$ values being: $T_{\text{final}}=0.5$ and $T_{\text{final}}=1.0$. We denote these two trajectories as $\text{Traj}_{\text{KA}}^{(1)}$ and $\text{Traj}_{\text{KA}}^{(2)}$, respectively.
III.3 Sample Labelling
Sample labelling is the assignment of a state set to its corresponding sample. In our study, we focus on the influence of particle rearrangement tendencies on the global phase structure. Thus, the state set of individual particles is {soft, hard}, while the global system exhibits a state set of {liquid, crystal, amorphous}.
The rationale behind this approach lies in the effectiveness of selecting {soft, hard} as the particle state set (or target property) in capturing the fundamental microscopic characteristics of particles. Previous studies [86, 81, 98] have demonstrated that machine learning models trained on particle rearrangement properties (i.e., soft/hard) as target variables are highly effective in exploring structural rearrangements [86, 81], dense packing structures [86, 81], non-exponential relaxation [81], glassy dynamics [81], and crystallization [98]. Representing these intrinsic properties (i.e., soft/hard) as a scalar field offers a robust framework for tracking their temporal evolution and serves as a powerful tool for investigating collective particle behavior at the system level [86, 81]. In contrast, defining particle states solely based on structural properties or symmetry, rather than employing the set of {soft, hard}, fails to accurately capture particle activity and, as a result, limits the ability to effectively study macroscopic phenomena [86, 81]. Therefore, setting the particle state set as {soft, hard} and the global system state set as {liquid, crystal, amorphous} must be done simultaneously.
We use MD simulations to obtain the data required for our research. Naturally, the data is stored in the form of simulation trajectories. The smallest accessible unit in a trajectory is a frame, and within each frame, the smallest accessible unit is the position and velocity vectors of an individual particle in the system at the corresponding time of this frame.
For system samples, we use MSD and $Q_{6}$ for phase labeling, denoting the label for trajectory $\text{Traj}(T)$ at time step $t$ as $y^{(T)}_{\text{entire},t}\in\{\text{liquid},\text{crystal},\text{amorphous}\}$. All data are listed in Table 7 of Appendix E for verification.
III.4 Dataset Construction
MD Protocol Protocol 1 Protocol 2
Group Group 1 Group 2 Group 3
Set Set 1 Set 3 Set 2 Set 4 Not applicable
Object Particle System Particle System Not applicable
Feature matrix $\mathbb{X}_{1}$ $\mathbb{X}_{3}$ $\mathbb{X}_{2}$ $\mathbb{X}_{4}$ Not applicable
Scale Local Global Local Global Not applicable
Property and/or status Rearrangement Tendency Global Phase Structure Rearrangement Tendency Global Phase Structure Not applicable
Status collection Soft or hard Crystal, amorphous or liquid Soft or hard Crystal, amorphous or liquid Amorphous or liquid
ML task Training Testing Not applicable
Task 1 Task 2 Task 1 Task 2 Not applicable
Not applicable Not applicable Global Softness Not applicable Global Softness
Fig. 10a) and 10b) Fig. 10c) and 10d)
Non-ML task Separation Index Separation Index Separation Index Separation Index Separation Index
Fig. 5a) and 6a) Fig. 5b) and 6b) Fig. 5a) and 6a) Fig. 5b), 6b), 10a) and 10b) Fig. 10c) and 10d)
System LJ system KA system
Dynamics Isothermal evolution Linear quenching
Potential Structural Transition (1) Crystallization from liquids, (2) Crystallization from amorphous solids, (3) Incomplete crystallization from liquids to amorphous solids, or (4) Remaining liquid phases (1) Vitrification, or (2) Remaining liquid phases
Table 1: Summary of MD simulation protocols, system categories, simulation dynamics, potential structural transitions, group classifications, extracted datasets, state sets, corresponding tasks, and the figures generated from these datasets. Note that “Not applicable” means that the content has no substantive meaning or is not relevant.
IV Experiments and Results
In this article, we conducted the following three computational experiments:
1. Applying PH descriptors to interpretable classification tasks: We demonstrated that PH-based descriptors are effective for interpretable classification tasks using SVMs. Moreover, we showed that a single variable is insufficient for soft-hard classification of particles but can successfully achieve the three-phase classification of the system. (Sec. IV.1).
2. Analyzing the Separation Index (SI): We computed the SI values and observed that as the neighborhood radius increases, the SI distribution of soft and hard particles gradually approaches that of liquid and solid systems. Moreover, the SI of global systems can serve as an approximate surrogate for explicitly mapping a single variable to global phase structures. (Sec. IV.2).
3. Tracking the temporal evolution of four metrics: We analyzed their variations along trajectories and compared them with $Q_{6}$ and MSD (Sec. IV.3).
IV.1 Machine Learning (ML)
IV.1.1 Task Description
Therefore, we conducted two classification tasks. Task 1 assesses the local descriptor $\mathbb{I}^{(P)}_{p}$ (Eq. 2) to determine whether a particle has a tendency for rearrangement (soft or hard), while Task 2 examines the global descriptor $\mathbb{I}^{(P)}_{\text{entire}}$ (Eq. 3) to classify phases throughout the entire system (liquid, crystal, or amorphous). Note that the relationship between data generation, dataset construction, and ML tasks is listed in Table 1.
The ML process follows the steps outlined below and is applicable to both Task 1 and Task 2:
1. We constructed four datasets, namely Set 1, Set 2, Set 3, and Set 4, according to the description in Sec. III.4 (or the summary provided in Table2). Each set corresponds to its respective feature matrix, denoted as $\mathbb{X}_{1}$, $\mathbb{X}_{2}$, $\mathbb{X}_{3}$, and $\mathbb{X}_{4}$.
2. We performed PCA on $\mathbf{X_{1}}$, $\mathbf{X}_{2}$, $\mathbf{X}_{3}$, and $\mathbf{X}_{4}$, ensuring that the explained variance ratio is $z$, and obtain the corresponding dimension-reduced matrices. Note that $\mathbf{X}_{2}$ should undergo the same PCA transformation as $\mathbf{X}_{1}$, while $\mathbf{X}_{4}$ should undergo the same PCA transformation as $\mathbf{X}_{3}$, as required in Sec. II.3.2. For convenience, we retain the original notation for the matrices after PCA transformation.
3. We trained our model using SVM, with its decision function uniformly expressed as $\mathbb{y}=f(\mathbb{X})$, evaluating both linear [82] and RBF kernels [99]. Here, $\mathbb{X}$ denotes the feature matrix corresponding to the training sets, i.e., $\mathbf{X}_{1}$ or $\mathbf{X}_{3}$, while $\mathbb{y}$ represents the label vector formed by concatenating $y^{(T)}_{p,t}$ (Task 1) or $y^{(T)}_{\text{entire},t}$ (Task 2) for each sample in the respective set. For Task 1, we trained the model on Set 1 and tested it on Set 2, while for Task 2, we trained on Set 3 and tested on Set 4.
Through this process, we not only derived the classification mappings for Task 1 and Task 2, $f:\mathbf{X}\to\mathbf{y}$, but also obtained the corresponding performance metrics, as reported in Table 2.
IV.1.2 ML Results and Analysis
The ML results are presented in Table 2. As described in Appendix A.1, the feature matrix $\mathbb{X}$ initially contains 1,600 columns, corresponding to a $\mathfrak{m}\times\mathfrak{n}$ grid with $\mathfrak{m}=\mathfrak{n}=40$, resulting in 1,600 predictors. However, applying PCA with an explained variance ratio of $z=98\%$ reduces the number of predictors to 188 for Task 1 and 122 for Task 2. Regardless of the kernel function used, Task 1 consistently achieves an accuracy above 86%, while Task 2 exceeds 95%. Overall, the linear kernel performs slightly better than the RBF kernel, though the difference remains marginal.